OpenStudy (jjuden):

Which relationship shows an inverse variation? x: 2,3,4,5 f(x):1,4,9,16 x:1,2,3,4 f(x):2,8,18,32 x:1,2,3,4 f(x):4,3,2,1 X:2,4,6,8 f(x):12,6,4,3

OpenStudy (anonymous):

I think inverse variation is 1/x

OpenStudy (anonymous):

c*(1/x)

OpenStudy (anonymous):

Just like last time plug in your numbers lets try first one x: 2,3,4,5 f(x):1,4,9,16 btw c*(1/x) = c/x :) same thing c/2=1 Multiply both sides by 2 you get c=2 ---- now we have formula 2/x =f(x) and we plug in 3 4 and 5 should be 4 9 16 2/3 right away t's not A :(

OpenStudy (jjuden):

ok i got you so far

OpenStudy (anonymous):

Lets do B x:1,2,3,4 f(x):2,8,18,32 c/x=y btw y=f(x) so c/x=f(x) same thing c/1=2 multiply by 1 c=2 2/x = f(x) f(2)=2/2=1 not 8 so it's not B either

OpenStudy (jjuden):

ok i follow so far i think its c am i right

OpenStudy (anonymous):

now for C x:1,2,3,4 f(x):4,3,2,1 c/x=y c/1=4 again just multiply by your current x :) c=4 f(2)=4/2=2 not 3 so it is not C either should be D but lets double check to make sure.

OpenStudy (anonymous):

X:2,4,6,8 f(x):12,6,4,3 c/x=y c/2=12 c=24 f(x)=24/x f(4)=24/4=6 checks out and so should the rest :D it's D